OpenStudy (anonymous):

Which of the following equations is the translation 2 units down of the graph of y = |x|? y = |x - 2| y = |x + 2| y = |x| - 2 y = |x| + 2

OpenStudy (anonymous):

y=\[\left| x \right|\]-2

OpenStudy (anonymous):

y = |x - 2| y = |x + 2| y = |x| - 2 y = |x| + 2